OpenStudy (anonymous):

How do write 125^2/3 in radical form? The choices are: A.15, 625 B. 125 C.20 D.25

OpenStudy (jdoe0001):

\(\Large \bf a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}}\)

OpenStudy (imstuck):

That looks like this:

OpenStudy (imstuck):

\[125^{\frac{ 2 }{ 3 }}\]The denominator of the exponent is always the index, like @jdoe0001 showed you. Get it?

OpenStudy (jdoe0001):

keep in mind that \(\bf 125\to 5\cdot 5\cdot 5\to 5^3\)

OpenStudy (anonymous):

yes that helps thank you both

OpenStudy (jdoe0001):

yw